Stock Valuation:               Primrose Corporation is considering investing in Metro Corporation by purchasing 10,000 shares...

  1. Stock Valuation:

      

       Primrose Corporation is considering investing in Metro Corporation by purchasing 10,000 shares of the corporation’s stock. The Metro Corporation just paid a dividend of $1.15. The dividends are expected to grow at 10 percent over the next 5 years. The company has a payout ratio of 40 percent and an expected PE ratio of 21. Based on the company’s payout ratio, earnings per share in year 5 would be $4.64 (EPS5 = D5 / Payout ratio = $1.86 / .40 = $4.64).

  1. What is the future (terminal value) stock price in five years? Show your work.

      

  1. What is the stock price today assuming a required return of 11 percent on this stock? Show your work.

      

  1. Assuming Primrose sells all $10,000 shares of stock at the end of the 5th year, how much money will the company gain or lose? Show your work.
Answer #1

A)

Given Expected PE ratio = 21 times

EPS in 5 years =$ 4.64

Terminal Value of stock Price in 5 years = 21 * 4.64 = $ 97.44

B)

Using dividend discount model,

Los 1.302 1.15 1.392 2.531 1.531 1.684 1.85 Primrose Corporation Dividends EPS in 5th year Terminal Value Net Cash flows Requ

Price of stock today = $ 63.42

C)

Profit per stock = (97.44 - 63.42) = $ 34.02

Total money gained for 10,000 shares = 34.02 * 10,000 = $ 340,200

Formulae

0.1 1.15 =C1071*(1+$B$1071) =D1071*(1+$B$1071) =E1071*(1+$B$1071) =F1071*(1+$B$1071) 1070 Primrose Corporation 1071 Dividends